I'm using code blocks to perform this. These problems are from my C Programming Beginners book. Not too familiar with

C programming so any help would be great. I would like to see how it is formatted. Also, I don't understand Problem 4 at all. Thanks.

1- Create three integer variables and initialize to zero.

2- Create three float variables and initialize to zero.

3- Create three double variables and initialize to zero.

4- Create one character variables and initialize to the space character.

5- Prompt the user for integer input and store in the first integer variable.

6- Prompt the user for integer input and store in the second integer variable.

7- Add the values stored in the two integers and assign it to the third integer.

8- Output the value of the third integer as shown in sample run.

9- Repeat steps 7 and 8 for subtraction, multiplication, and division.

10- Repeat steps 5 through 9 for floats and doubles accordingly. Output float with precision

of 2 and doubles with precision of 4.
Answer #1

ANSWER:

question 1 to 4 is quite straight forward , just declare a variable and assign 0 to it .Which you can easily understand ow to do it by looking at the code thoroughly once, then in question 5 and 6 we, have to prompt the user for entering new value for our declared variables of various data type one by one, so in c for user input , we use scanf() function, in it we use different format specifier for different datatype of input, for integer we have used %d for float we have used %f and for double we have used %lf.

then we have to do binary operations like addition , subtraction, multiplication and division and store the value in the third remaining variable.and then print it using the right format specifier as shown in the code and  as used during the use of scanf() function.

In question 8, you are asking about the output as in the sample run, but yu have not provided any sample run screenshot, so i have printed the output in my own way , if it's different than yours, please comment what was in the sample run, i'll help you then.

Question 9 is just the repetition of the above steps, just for different datatype , so we have done accordingly,

Question 10 we have to print the float result with precision value to 2, means after decimal we want only 2 digits , and precision value to 4 in case of double, means 4 digits after the decimal value in the result, so we can set precision of decimal values by using %0.2f in case of float and %0.4lf in case of double here.

You can get more clarity when you go through the code , i have inserted enough comments to help you understand each line.

NOTE: I have used online c compiler.

Code:

#include <stdio.h>

int main()
/*question number 1*/
{ int x; // creating first integer variable
int y; // creating second integer variable
int z; // creating third integer variable
x = 0; // initialising x to zero(first integer)
y = 0; // initialising y to zero(second integer)
z = 0; // initialising z to zero(third integer)
  
/*question number 2*/
float p; // creating first float variable
float q; // creating second float variable
float r; // creating third float variable
p = 0; // initialising p to zero(first float)
q = 0; // initialising q to zero(second float)
r = 0; // initialising r to zero(third float)
  
/*question number 3*/
double a; // creating first double variable
double b; // creating second double variable
double c; // creating third double variable
a = 0; // initialising a to zero(first double)
b = 0; // initialising b to zero(second double)
c = 0; // initialising c to zero(third double)
  
/*question number 4*/
char ch; // creating char variable named ch
ch = " "; // initialising char ch to space character
  
/*question number 5*/
printf("Enter the new value of first integer :"); // printing message for prompting
scanf("%d",&x); // taking user input using scanf
  
/*question number 6*/
printf("Enter the new value of second integer :"); // printing message for prompting
scanf("%d",&y); // taking user input using scanf
  
/*question number 7*/
z = x + y; // adding variable x and y and storing the sum into z
  
/*question number 8*/
printf("sum of %d and %d is equal to %d \n",x,y,z); // output the value of third variable z
  
/*question number 9*/
z = x - y; // subtracting y from x and storing the result in z
printf("Difference of %d and %d is equal to %d \n",x,y,z); // output the value subtraction
  
z = x * y; // multiplication of x and y storing result in z
printf("Product of %d and %d is equal to %d \n",x,y,z); // output the Product value
  
z = x / y; // division of a and y storing result in z
printf("Division of %d and %d is equal to %d \n\n",x,y,z); // output the integer quotient value

  
/*question number 10*/
  
// **************************for float***********************
  
/*repeating question 5*/
printf("Enter the new value of first float :"); // printing message for prompting
scanf("%f",&p); // taking user input using scanf
  
/*repeating question 6*/
printf("Enter the new value of second integer :"); // printing message for prompting
scanf("%f",&q); // taking user input using scanf
  
/*repeating question 7*/
r = p + q; // adding variable p and q and storing the sum into r
  
/*question number 8*/
printf("sum of %f and %f is equal to %0.2f \n",p,q,r); // output the value of third variable r
  
/*question number 9*/
r = p - q; // subtracting q from p and storing the result in r
printf("Difference of %f and %f is equal to %0.2f \n",p,q,r); // output the subtraction value(using %0.2f for precision upto 2 digit after decimal)
  
r = p * q; // multiplication of p and q storing result in r
printf("Product of %f and %f is equal to %0.2f \n",p,q,r); // output the Product value(using %0.2f for precision upto 2 digit after decimal)
  
r = p / q; // division of p and q storing result in r
printf("Division of %f and %f is equal to %0.2f \n\n",p,q,r); // output the integer quotient (using %0.2f for precision upto 2 digit after decimal)
  
  
  
// ***********************for double***************************
  
  
  
/*repeating question 5*/
printf("Enter the new value of first double :"); // printing message for prompting
scanf("%lf",&a); // taking user input using scanf
  
/*repeating question 6*/
printf("Enter the new value of second double :"); // printing message for prompting
scanf("%lf",&b); // taking user input using scanf
  
/*repeating question 7*/
c = a + b; // adding variable a and b and storing the sum into c
  
/*question number 8*/
printf("sum of %lf and %lf is equal to %0.4lf \n",a,b,c); // output the value of third variable c
  
/*question number 9*/
c = a - b; // subtracting q from p and storing the result in r
printf("Difference of %lf and %lf is equal to %0.4lf \n",a,b,c); // output the subtraction value(using %0.4f for precision upto 4 digit after decimal)
  
c = a * b; // multiplication of p and q storing result in r
printf("Product of %lf and %lf is equal to %0.4lf \n",a,b,c); // output the Product value(using %0.4f for precision upto 4 digit after decimal)
  
a = b / c; // division of p and q storing result in r
printf("Division of %lf and %lf is equal to %0.4lf \n",a,b,c); // output the integer quotient (using %0.4f for precision upto 4 digit after decimal)

  
return 0;
}

OUTPUT: